OpenStudy (anonymous):

If X varies directly as Y, and X=35 when Y=5, find X when Y=7.

OpenStudy (alexwee123):

Y=K/X 5=k/35 k=175

OpenStudy (alexwee123):

y=kx 7=175/x 7x=175 x=25

OpenStudy (anonymous):

how sure are you? my grade depends on this!

OpenStudy (alexwee123):

yes i'm very sure :) you need an explanation?

OpenStudy (anonymous):

please

OpenStudy (alexwee123):

well inverse variation equation is xy=k

OpenStudy (alexwee123):

where k is the constant of variation

OpenStudy (alexwee123):

so i used the equation to find the "k" by plugging in x=35 and y=5 and k=175 then i plugged k back into the equation and as well as y, (which =7), and found out the value for x

OpenStudy (alexwee123):

and i tweaked the equation from xy=k to y=k/x because i'm most comfortable with it but you don't have to use it that way even though both would give the same values